# It’s complicated. The relationship of algorithmic fairness and non-discrimination regulations in the EU AI Act

### v1

@@ Abstract.
What constitutes a fair decision? This question is not only difficult for humans but becomes more challenging when Artificial Intelligence (AI) models are used. In light of discriminatory algorithmic behaviors, the EU has recently passed the AI Act, which mandates specific rules for AI models, incorporating both traditional legal non-discrimination regulations and machine learning based algorithmic fairness concepts. This paper aims to bridge these two different concepts in the AI Act through: First a high-level introduction of both concepts targeting legal and computer science-oriented scholars, and second an in-depth analysis of the AI Act’s relationship between legal non-discrimination regulations and algorithmic fairness. Our analysis reveals three key findings: (1.), most non-discrimination regulations target only high-risk AI systems. (2.), the regulation of high-risk systems encompasses both data input requirements and output monitoring, though these regulations are often inconsistent and raise questions of computational feasibility. (3.) Regulations for General Purpose AI Models, such as Large Language Models that are not simultaneously classified as high-risk systems, currently lack specificity compared to other regulations. Based on these findings, we recommend developing more specific auditing and testing methodologies for AI systems. This paper aims to serve as a foundation for future interdisciplinary collaboration between legal scholars and computer science-oriented machine learning researchers studying discrimination in AI systems.
@@ 2.1. Algorithmic fairness: A gentle introduction
Direct discrimination is defined as situations in which “one person is treated less favourably than another is, has been or would be treated in a comparable situation” (Article 2(2)(a) Directive 2000/43/EC). This means that individuals are discriminated against (directly) by using a (protected) attribute. In algorithmic contexts, an example would be a simple algorithm that grants anyone below 50 years a loan and denies it to anyone above 50 years. In this setting, only the protected attribute age is used to define the loan threshold. People would be directly discriminated against based on their age.
Indirect discrimination is much more challenging to tackle. In indirect discrimination cases, seemingly neutral attributes are used, but they rely on a protected attribute. An example of indirect discrimination would be if not the age is used but the amount of gray hair, which is usually highly correlated with age. If an algorithm does not use age directly but the amount of gray hair, then people are indirectly discriminated against based on their age.
Also, a third category of discrimination needs to be taken into account which challenges traditional non-discrimination law: intersectional discrimination. Intersectional discrimination concerns cases “originating in several inextricably linked vectors of disadvantage” (Xenidis, 2023). This becomes especially important since (modern) AI models do not use single variables as input — such as our simplified loan example based on the single variable above — but instead use many different aspects as input. This can lead to effects where discrimination may occur at the intersection of gender and age. Importantly, if you only investigate age and gender separately, no discrimination would be recognized (Weerts et al., 2023). It is an open question if and how, at the current stage, the European Court of Justice recognizes intersectional discrimination as a form of discrimination (Weerts et al., 2023; Xenidis, 2023; Atrey, 2018).
With these three types of discrimination based on EU discrimination law, the last concept that needs to be discussed is the concept of Fundamental Rights which is mentioned in multiple articles in the AI Act.
@@ 56 Codes of practice for GPAI Models
What is a high-risk AI system? Article 6 AIA defines different aspects of high risks. In combination with Annex III, high-risk systems are, for example, intended to be used as safety components in the management and operation of critical digital infrastructure (Annex III(2) AIA) or systems intended to be used for the recruitment or selection of natural persons (Annex III(4)(a) AIA). The reasoning behind the specific regulations in Article 6 AIA is that systems which pose a specific risk to fundamental rights or safety need tighter regulation (Fraser and y Villarino, 2024) than other less risky systems. It is expected that 5%-15% [^9] For a critical assessment of these numbers see (Almada and Petit, 2023). of all AI systems fall under the category of high-risk systems (Commission, 2021).
We identified Articles 9, 10, and 15 AIA as main non-discrimination provisions of high-risk systems. These will be discussed in detail in the following paragraphs [^10] Please note that we do not differentiate here between the obligations for providers and those obligations for the deployers of AI systems, since we focus on the interaction of computer science and law.. All other Articles listed in tab 1 will be discussed and summarized in the last paragraph.
@@
Important in the context of non-discrimination law is Article 9(2)(a) AIA, which requires the identification and analysis of known and reasonably foreseeable risks that the high-risk AI system can pose to health, safety, or Fundamental Rights. As explained in Section 2.2, Fundamental Rights also include the right of non-discrimination. Thus, indirectly, Article 9 AIA already requires the identification and analysis of non-discrimination issues. Zuiderveen Borgesius et al. (2024) pointed out that this Article mandates the identification and mitigation of discrimination in AI systems. The initial draft of the AI Act did not include the reference to Fundamental Rights, and it was added by the EU Council in 2022 111111see Council of the European Union, “Proposal for a Regulation of the European Parliament and the Council laying down harmonized rules on artificial intelligence (Artificial Intelligence Act) and amending certain Union legislative acts – General approach” Doc. 14954/22, 25 November 2022. (Schuett, 2024), highlighting the need to ensure non-discriminatory regulations.
@@
Second, the link to Fundamental Rights leaves many potential non-discrimination factors unaddressed (Kusche, 2024). According to Article 21(1) CFR, discrimination based on sex, race, color, ethnic or social origin, genetic features, language, religion or belief, political or any other opinion, membership of a national minority, property, birth, disability, age, or sexual orientation shall be prohibited. However, discrimination based on other factors is not prohibited. This creates unnecessary opportunities for discrimination. Furthermore, as explained above, the Fundamental Rights need to be balanced; the right of non-discrimination is only one among others. It is therefore unclear how effective this protection is (Kusche, 2024).
@@
Input bias data detection, prevention, and mitigation methods remain largely unexplored. These methods are not clearly defined in the AI Act (Wachter, 2024; Deck et al., 2024b). Article 10(2)(g) AIA appears to be inspired by the technical literature on fairness metrics for algorithmic outputs. However, compared to output fairness metrics, technical bias detection methods for input data are understudied. Furthermore, as previously noted, the causes of biases in ML are multi-causal (Gebru, 2019; Gebru et al., 2021; Hanna et al., 2020; Mehrabi et al., 2021; Bender et al., 2021). Even when data are collected representatively, biases can occur (see (Mehrabi et al., 2021) for a comprehensive list). Two examples illustrate this argument:
@@
The data governance strategies in Article 10(2)(g) AIA focus too heavily on technical aspects of input data biases. While systematic fairness approaches have been developed for output bias detection, prevention, and mitigation (Han et al., 2023; Krasanakis and Papadopoulos, 2024; Pagano et al., 2023), we lack similarly comprehensive approaches for input data analysis. This creates a significant gap between the legal requirements in Article 10(2)(g) AIA and their technical and ethical feasibility.
@@
Summary: High-risk systems show vague non-discrimination regulations. Regulations regarding high-risk systems are vague. In contrast to the ambitions of the AI Act, the concrete non-discrimination regulations lack specificity. For the ML community it is challenging to understand the implications of the AI Act. Only Article 10 AIA and Article 15 AIA explicitly mention non-discrimination and biases, compared to the vague statement of Fundamental Rights protection in Article 9 AIA. However, Article 10 AIA only considers the input phase, while Article 15 AIA only addresses the output phase for feedback loops. It would be beneficial if Article 15 AIA, Article 9 AIA, and Article 10 AIA regulated the same types of direct, indirect, and intersectional discrimination on both the input and output sides.
@@ 3.4. General Purpose AI non-discrimination regulations
Summary: GPAI models are only vaguely regulated. In summary, GPAI models are currently only vaguely regulated. There exists significant regulatory uncertainty, especially regarding how the regulations of the AI Act can be implemented in practice. Perhaps the Code of Practice being developed by the AI Office and the European Commission can provide guidance in this area.
@@ 4. Summary & Outlook: How to proceed?
As Mayson (2018) already pointed out: Bias in, bias out. The AI act seek to complement existing non-discrimination EU law with it processed based regulation (Zuiderveen Borgesius et al., 2024). However, even on this broader level, the AI Act fails to account for major issues.
@@
Second, the AI Act does not include any regulations for none to limited-risk systems. Many AI systems are not regulated at all within the AI Act. It is worth noting that the EU estimated that high-risk systems comprise roughly 5%-15% of all AI systems in the EU. Although the remaining 85%-95% of all systems include a few GPAI systems, most AI systems remain entirely uncovered. Non-model specific regulations, such as the obligation to foster AI literacy in Article 4 AIA, are insufficient to ensure non-discriminatory practices. From our perspective, the legislator is missing an opportunity to regulate AI systems on a broader scope. This concerns not only the obligations but also specific rights. For example, the use of personal data for bias detection could also be beneficial for other than high-risk systems.
@@
Many of the rules focus on (internal) compliance (Zuiderveen Borgesius et al., 2024) and checklist rather than on output-based safeguards for users affected by the AI system This internal compliance is also linked to the self-assessment approach. It is questionable whether the self-assessment is suitable for protecting individuals and their Fundamental Rights as well as non-discrimination regulation (Bosoer et al., 2023).
@@
Future: A call for more interdisciplinary work. While the AI Act is casuistic and incomplete at the same time (Ződi, 2024), we believe that the AI Act would benefit from four refinements. (1.) If the legislators decide to target non-discrimination regulation, non-discrimination should be explicitly mentioned. In contrast to vague terms such as bias, Fundamental Rights, or fairness, non-discrimination is a well-established legal concept within other European legislative frameworks. (2.) The non-discrimination regulation should cover input as well as output parts of the algorithms in one single article, also along the full AI life cycle (Deck et al., 2024b). Algorithmic output in terms of non-discrimination is only regulated if systems continuously learn after being placed on the market. Otherwise, outputs in terms of discriminatory results of AI systems are not regulated at all. Technical and non-technical means for bias detection, prevention, and mitigation for input and output data should be developed. (3.) Clarification is needed on appropriate measures for non-discrimination. While a reverse burden-of-proof approach is possible here (Wachter, 2024), similar processes as the Code of Practice for GPAI models might be possible such that technical and legal experts as well as all stakeholders work together on refined non-discrimination measures. (4.) If GPAI models are regulated in the current form, a closer link between GPAI models and downstream applications needs to be established (Hacker et al., 2024b). Responsibilities for non-discrimination law compliance must be specified. It is especially important to note that this work is best done in joint community efforts of legal researchers and ML researchers. A bridge between these two fields is highly desirable since regulation should be guided by technical appropriateness and feasibility.
@@ 5. Acknowledgments
We would like to thank Miriam Rateike for her detailed comments on an earlier version of the manuscript. We are grateful for comments from Christoph Kern and the Social Data Science and Statistical Learning group during an institute seminar, which clearly improved this paper. Furthermore, we thank Alina Wenick, Tharushi Abeynayaka and Mara Seyfert for comments on the manuscript and proofreading the paper.

### v2

@@ Abstract.
What constitutes a fair decision? This question is not only difficult for humans but becomes more challenging when Artificial Intelligence (AI) models are used. In light of discriminatory algorithmic behaviors, the EU has recently passed the AI Act, which mandates specific rules for AI models, incorporating both traditional legal non-discrimination regulations and machine learning based algorithmic fairness concepts. This paper aims to bridge these two different concepts in the AI Act through: First a high-level introduction of both concepts targeting legal and computer science-oriented scholars, and second an in-depth analysis of the AI Act’s relationship between legal non-discrimination regulations and algorithmic fairness. Our analysis reveals three key findings: (1.), most non-discrimination regulations target only high-risk AI systems. (2.), the regulation of high-risk systems encompasses both data input requirements and output monitoring, though these regulations are partly inconsistent and raise questions of computational feasibility. (3.) Regulations for General Purpose AI Models, such as Large Language Models that are not simultaneously classified as high-risk systems, currently lack specificity compared to other regulations. Based on these findings, we recommend developing more specific auditing and testing methodologies for AI systems. This paper aims to serve as a foundation for future interdisciplinary collaboration between legal scholars and computer science-oriented machine learning researchers studying discrimination in AI systems.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 2025 ACM Conference on Fairness, Accountability, and Transparency; ;
@@ 2.1. Algorithmic fairness: A gentle introduction
Direct discrimination is defined as situations in which “one person is treated less favourably than another is, has been or would be treated in a comparable situation” (Article 2(2)(a) Directive 2000/43/EC). This means that individuals are discriminated against by using a (listed) protected attribute. In algorithmic contexts, an example would be a simple algorithm that grants anyone below 50 years a loan and denies it to anyone above 50 years. In this setting, only the protected attribute age is used to define the loan threshold. People would be directly discriminated against based on their age.
Indirect discrimination is much more challenging to tackle. In indirect discrimination cases, seemingly neutral attributes are used, but they rely on a protected attribute. An example of indirect discrimination would be if not the age is used but the amount of gray hair, which is usually highly correlated with age. If an algorithm does not use age directly but the amount of gray hair, then people are indirectly discriminated against based on their age. For indirect discrimination, it is important to note that it can be justified through legitimate aims and appropriate means (Zuiderveen Borgesius et al., 2024).
Also, a third category of discrimination needs to be taken into account [^9] Please note that other non-discrimination rules in sector-specific regulations, such as the EU Consumer Credit Directive, are beyond the scope of our paper., which challenges traditional non-discrimination law: intersectional discrimination. Intersectional discrimination concerns cases “originating in several inextricably linked vectors of disadvantage” (Xenidis, 2023). This becomes especially important since (modern) AI models do not use single variables as input — such as our simplified loan example based on the single variable above — but instead use many different aspects as input. This can lead to effects where discrimination may occur at the intersection of gender and age. Importantly, if you only investigate age and gender separately, no discrimination would be recognized (Weerts et al., 2023). It is an open question if and how, at the current stage, the European Court of Justice recognizes intersectional discrimination as a form of discrimination (Weerts et al., 2023; Xenidis, 2023; Atrey, 2018).
With these three types of discrimination based on non-discrimination law, the last concept that needs to be discussed is the concept of Fundamental Rights which is mentioned in multiple articles in the AI Act.
@@
Further complications arise due to the fact that the provisions of the CFR target public parties in relation to private parties (vertical applicability) (Fornasier, 2015). Therefore, there are debates about how exactly the CFR applies to relationships between private parties (horizontal applicability) in general (Fornasier, 2015; Cherednychenko, 2007; Frantziou, 2015; Prechal, 2020) and in the AI Act specifically (Lewis et al., 2025).
@@ 3.1. The EU AI Act: History, Scope, and Definitions
Providers (Article 3(3) AIA) and deployers (Article 3(4) AIA) are the central addressees of the AI Act. Providers are the entities that develop and place a model on the market or put the system into service. In contrast, the deployers are using an AI system under its authority. Many of the following obligations focus on the providers of a system (Article 16 AIA) compared to the deployers (Article 26 AIA).
@@ 56 Codes of practice for GPAI Models
What is a high-risk AI system? Article 6 AIA defines different aspects of high risks. In combination with Annex III, high-risk systems are, for example, intended to be used as safety components in the management and operation of critical digital infrastructure (Annex III(2) AIA) or systems intended to be used for the recruitment or selection of natural persons (Annex III(4)(a) AIA). The reasoning behind the specific regulations in Article 6 AIA is that systems which pose a specific risk to fundamental rights or safety need tighter regulation (Fraser and y Villarino, 2024) than other less risky systems. It is expected that 5%-15% [^10] For a critical assessment of these numbers see (Almada and Petit, 2023) of all AI systems fall under the category of high-risk systems (Commission, 2021).
We identified Articles 9, 10, and 15 AIA as main non-discrimination provisions of high-risk systems. These will be discussed in detail in the following paragraphs [^11] Please note that we do not differentiate here between the obligations for providers and those obligations for the deployers of AI systems, since we focus on the interaction of computer science and law. All other Articles listed in tab 1 will be discussed and summarized in the last paragraph.
@@
Important in the context of non-discrimination law is Article 9(2)(a) AIA, which requires the identification and analysis of known and reasonably foreseeable risks that the high-risk AI system can pose to health, safety, or Fundamental Rights. As explained in Section 2.2, Fundamental Rights also include the right of non-discrimination. Thus, indirectly, Article 9 AIA already requires the identification and analysis of non-discrimination issues. Zuiderveen Borgesius et al. (2024) pointed out that this Article mandates the identification and mitigation of discrimination in AI systems. The initial draft of the AI Act did not include the reference to Fundamental Rights, and it was added by the EU Council in 2022 121212see Council of the European Union, “Proposal for a Regulation of the European Parliament and the Council laying down harmonized rules on artificial intelligence (Artificial Intelligence Act) and amending certain Union legislative acts – General approach” Doc. 14954/22, 25 November 2022. (Schuett, 2024), highlighting the need to ensure non-discriminatory regulations.
@@
Second, the link to Fundamental Rights leaves many potential non-discrimination factors unaddressed (Kusche, 2024). According to Article 21(1) CFR, discrimination based on sex, race, color, ethnic or social origin, genetic features, language, religion or belief, political or any other opinion, membership of a national minority, property, birth, disability, age, or sexual orientation shall be prohibited. However, discrimination based on other factors is not prohibited by Article 21(1) CFR but may be prohibited by other European or national laws. This creates unnecessary opportunities for discrimination. Furthermore, as explained above, the Fundamental Rights need to be balanced; the right of non-discrimination is only one among others. It is therefore unclear how effective this protection is (Kusche, 2024).
@@
Input bias data detection, prevention, and mitigation methods remain largely unexplored. These methods are not clearly defined in the AI Act (Wachter, 2024; Deck et al., 2024b). Article 10(2)(g) AIA appears to be inspired by the technical literature on fairness metrics for algorithmic outputs. Nevertheless, Article 10(2)(g) AIA targets the input of algorithms. Therefore, classical fairness metrics for algorithmic outputs do not apply.
However, compared to output fairness metrics, technical bias detection methods for input data are understudied. Furthermore, as previously noted, the causes of biases in ML are multi-causal (Gebru, 2019; Gebru et al., 2021; Hanna et al., 2020; Mehrabi et al., 2021; Bender et al., 2021). Even when data are collected representatively, biases can occur (see (Mehrabi et al., 2021) for a comprehensive list). Two examples illustrate this argument:
@@
The data governance strategies in Article 10(2)(g) AIA focus too heavily on technical aspects of input data biases. While systematic fairness approaches have been developed for output bias detection, prevention, and mitigation (Han et al., 2023; Krasanakis and Papadopoulos, 2024; Pagano et al., 2023; Zollo et al., 2024), we lack similarly comprehensive approaches for input data analysis. This creates a significant gap between the legal requirements in Article 10(2)(g) AIA and their technical and ethical feasibility.
@@
Summary: High-risk systems balancing between specificity and generality The AI Act aims to be a specific law for AI. Classical EU non-discrimination law can already be difficult to interpret (see related literature above). In contrast to the ambitions of the AI Act, the concrete non-discrimination regulations lack specificity. For the ML community it is challenging to understand the implications of the AI Act. Only Article 10 AIA and Article 15 AIA explicitly mention non-discrimination and biases, compared to the vague statement of Fundamental Rights protection in Article 9 AIA. However, Article 10 AIA only considers the input phase, while Article 15 AIA only addresses the output phase for feedback loops. It would be beneficial if Article 15 AIA, Article 9 AIA, and Article 10 AIA regulated the same types of direct, indirect, and intersectional discrimination on both the input and output sides. In this case, the AI Act would substantially add specific benefits to classical broad-applicability non-discrimination law.
@@ 3.4. General Purpose AI non-discrimination regulations
Summary: GPAI models are only vaguely regulated. In summary, GPAI models are currently only vaguely regulated. There exists significant uncertainty, especially regarding how the regulations of the AI Act can be implemented in practice for GPAI models, besides attempts to solve this (Zollo et al., 2024). Perhaps the Code of Practice being developed by the AI Office and the European Commission can provide guidance in this area.
@@ 4. Summary & Outlook: How to proceed?
As Mayson (2018) already pointed out: Bias in, bias out. The AI act seek to complement existing non-discrimination law with it processed based regulation (Zuiderveen Borgesius et al., 2024). However, even on this broader level, the AI Act fails to account for major issues.
@@
Second, the AI Act does not include many regulations for none to limited-risk systems. The AI Act not only imposes obligations on high-risk systems but also grants rights, such as the authorization to use special categories of personal data (Article 9 GDPR) in Article 10(5) AIA in high-risk AI contexts. Such permission could also be granted for non-high-risk systems. Furthermore, Such permission could also be granted for non-high-risk systems. Furthermore, non-model specific regulations, such as the obligation to foster AI literacy in Article 4 AIA, are insufficient to ensure non-discriminatory practices. From our perspective, legislators missed an opportunity to regulate AI systems on a broader scope.
@@
Many of the rules focus on (internal) compliance (Zuiderveen Borgesius et al., 2024) and checklist rather than on output-based safeguards for users affected by the AI system. This internal compliance is also linked to the self-assessment approach. It is questionable whether the self-assessment is suitable for protecting individuals and their Fundamental Rights as well as non-discrimination regulation (Bosoer et al., 2023).
@@
Future: A call for more interdisciplinary work. While the AI Act is casuistic and incomplete at the same time (Ződi, 2024), we believe that the AI Act would benefit from four refinements. (1.) If the legislators decide to target non-discrimination regulation, non-discrimination should be explicitly mentioned. In contrast to terms such as bias or fairness, non-discrimination is a well-established legal concept within other European legislative frameworks. (2.) The non-discrimination regulation should cover input as well as output parts of the algorithms in one single article, also along the full AI life cycle (Deck et al., 2024b). Algorithmic output in terms of non-discrimination is only regulated if systems continuously learn after being placed on the market. Otherwise, outputs in terms of discriminatory results of AI systems are not regulated at all. Technical and non-technical means for bias detection, prevention, and mitigation for input and output data should be developed. (3.) Clarification is needed on appropriate measures for non-discrimination. While a reverse burden-of-proof approach is possible here (Wachter, 2024), similar processes as the Code of Practice for GPAI models might be possible such that technical and legal experts as well as all stakeholders work together on refined non-discrimination measures. (4.) If GPAI models are regulated in the current form, a closer link between GPAI models and downstream applications needs to be established (Hacker et al., 2024b). Responsibilities for non-discrimination law compliance must be specified. It is especially important to note that this work is best done in joint community efforts of legal researchers and ML researchers. A bridge between these two fields is highly desirable since regulation should be guided by technical appropriateness and feasibility.
@@ 5. Acknowledgments
We would like to thank Miriam Rateike for her detailed comments on an earlier version of the manuscript. We are grateful for comments from Christoph Kern and the Social Data Science and Statistical Learning group during an institute seminar, which clearly improved this paper. Furthermore, we thank Alina Wenick, Tharushi Abeynayaka, and Mara Seyfert for their comments on the manuscript and for proofreading the paper.